# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v1

@@ IV-G1 Knowledge Base Question Answering
ReTraCk [239] links entities using mention detection, and retrieves schema items using dense retriever Bert. It then generates logical forms by LSTM, incorporating retrieved items through knowledge-specific rules. For a given query, Unseen Entity Handling [51] retrieves topic entities through FreeBase [240], and concatenates the query with the entity for an encoder-decoder to generate SPARQL output. CBR-KBQA [52] retrieves relevant questions and corresponding logical form answers with roberta-based deep retrieval, then concatenates the question and the retrieved pairs for encoder-decoder transformer model. It also revises the final generation result to align the generated relations with relations present in the local neighborhood of the query entity in the knowledge graph. GMT-KBQA [50] first retrieves relevant entities and relations through bert-based deep retrieval, then concatenates the information for T5 generation. To improve the retrieval result, it leverages cross-encoder to rerank the candidates, and uses the same T5 structure to conduct relation classification and entity dismbiguation. RNG-KBQA [121] enumerates candidate logical forms in the knowledge graph, then ranks the candidates and concatenates them with query to form the prompt input to generate final logical form through a T5 model. Based on this idea, TIARA [122] also retrieve entity and schema besides logical forms, while a following work [241] retrieves top-k questions with BM25. Uni-Parser [133] retrieves relevant entities from knowledge graph using mention detection, cross-encoder ranker, and 2-hop paths extraction. It also considers enumerating tables and columns from databases. On obtaining the relevant information, it concatenates the top-k primitives with the query and generates logical forms through T5. UniK-QA [142] concatenates the text forms of the components in a triplet and build document pool for retrieval. For a given question, it leverages dense retriever for relevant documents, then performs fusion-in-decoder technique to incorporate the information for answer generation. SKP [144] performs similar operations as UniK-QA in inference, but adds a pretraining stage with a knowledge-aware MLM loss on triplets and knowledge constrastive loss with respect to the retrieved items. BLLM augmentation [135] uses TIARA as the retrieval for relevant knowledge base elements, then performs in-context learning through black-box LLM such as GPT-4 for generating logical forms. ECBRF [134] follows the case-based reasoning paradigm [242], retrieving similar triplet with dense retriever and constructing prompt input for BART or GPT-2 in-context learning. DECAF [143] forms Resource Description Format knowledge base triplets into sentences, then concatenates sentences with the same head entity into documents. It retrieves relevant documents using BM25 sparse retrieval or Bert dense retrieval, then leverages Fusion-in-Decoder technique to generate logical form and direct answer given each (query, document) pair as input. It combines the output to obtain the final answer. KD-CoT [145] uses the same dense retriever and fusion-in-decoder generator as DECAF. It follows a Chain-of-Thought paradigm, iteratively performing retrieval, generation, and verification until the CoT is finished. FC-KBQA [243] extracts relevant class, relation, and entity given a question. For class and relation, it uses BM25 as retriever and a Bert-based cross-encoder as re-ranker. For entity, it follows the mention detection paradigm. To compose all the component candidates, it applies T5 model to generate logical expression. Keqing [244] first decomposes a complext question into simple sub-questions through finetuned LLM, then retrieves similar sub-question template by dense retriever RoBERTa to extract candidate entities from knowledge graph, and finally generates answer through ChatGPT given relevant entities as context input. To probe the deep understanding of natural language in LLMs with formal languages, a research work [245] explores the capability of formal language understanding and formal language generation. It leverages retrieved pairs to perform in-context learning. For understanding, it uses tree edit distance to retrieve similar logical forms, while for generation, it uses BM25 to retrieve similar natural language queries.
IV-G2 Knowledge Graph Completion
Knowledge graph is consisted of triplets, including head entity, relation, and tail entity. The task of knowledge graph completion is to predict the missing entity in an incomplete triplet.
ReSKGC [146] linearizes all training triplets into text by concatenation, then retrieves relevant triplets using BM25, and generates completed triplet using T5 with fusion-in-decoder.
@@ IV-H RAG for Science
RAG has also emerged as a promising research direction for many interdisciplinary applications, such as strengthening molecular generation and guiding medical tasks.
@@ IV-H1 Drug Discovery
The goal of drug discovery is to generate molecules that concurrently fulfill diverse properties. RetMol [53] incorporates a lightweight retrieval mechanism into a pre-trained encoder-decoder generative model. This approach adopts the SMILES string [246] representation of molecules along with the ChemFormer model [247] , involving the retrieval and fusion of exemplar molecules with the input molecule. It also utilizes an iterative refinement process to dynamically update both the generated molecules and the retrieval database. Targeted at structure-based drug design (SBDD) [248, 249], PromptDiff [250] proposes an innovative interaction-based retrieval-augmented 3D molecular diffusion model to facilitate target-aware molecule generation. PromptDiff retrieves a curated set of ligand references, i.e., those with desired properties such as high binding affinity, to steer the diffusion model towards synthesizing ligands that satisfy design criteria. DeepICL [147] constitutes an interaction-focused 3D molecular generative framework with a specific emphasis on modeling protein-ligand interactions, aimed at facilitating a generalizable pocket-constrained ligand design.Utilizing a VAE architecture [251], the encoder module processes protein-ligand complexes into a latent vector, which is combined with retrieved prior knowledge from interaction conditions. This combined information guides the sequential, atom-wise generation of the ligand structure by the decoder module, ensuring that ligand atoms are appropriately placed to facilitate the desired interaction with the target.
IV-H2 Medical Applications
Several recent studies have improved the expressiveness of LLM by retrieving information from biomedical domain-specific databases, thereby augmenting the model’s capabilities to provide valuable guidance for tasks in the medical field. GeneGPT [54] is a novel method in training a LLM to effectively employ the Web API offered by the National Center for Biotechnology Information (NCBI) for addressing genomics inquiries. It enhances the LLM by integrating domain-specific tools, thereby refining its ability to access biomedical information through the retrieval of such data from the NCBI via API calls. When tasked with GeneTuring tests, GeneGPT instructs Codex [2] to utilize NCBI Web APIs through in-context learning and an augmented decoding algorithm proficient in detecting and executing API calls. Chat-Orthopedist [136] establishes an external knowledge base with information on adolescent idiopathic scoliosis (AIS) disease and treatment options, and develops a retrieval-augmented ChatGPT to feed LLMs with AIS domain knowledge, providing accurate and comprehensible responses to patient inquiries. Utilizing an external AIS knowledge base and a dense retriever, Chat-Orthopedist enhances LLMs with domain-specific knowledge related to scoliosis, avoiding the issue of “molecular hallucinations”.
@@ V Benchmark
CRUD-RAG [256] divides all RAG tasks into four categories, which are Create, Read, Update, and Delete, respectively, and also evaluates each category using text continuation, question answering (with single- and multi-document questions), hallucination modification, and open-domain multi-document summary.
@@ VI-B2 Efficient Deployment and Processing
Currently, several deployment solutions of query-based RAG for LLMs have been proposed, such as LangChain [263] and LLAMA-Index [169]. However, for other types of RAG and generation tasks, there lacks a plug-and-play solution. In addition, given the extra overhead introduced by retrieval, and considering that the complexities of both the retriever and generator will continue to grow, achieving efficient processing in RAG remains a challenge, necessitating targeted system optimization.

### v2

@@ IV-G1 Knowledge Base Question Answering
Query-based RAG is the mainstream approach. For a given query, Unseen Entity Handling [52] retrieves topic entities through FreeBase [321], and concatenates the query with the entity for an encoder-decoder to generate SPARQL output. CBR-KBQA [53] retrieves relevant questions and corresponding logical form answers with roberta-based deep retrieval, then concatenates the question and the retrieved pairs for encoder-decoder transformer model. It also revises the final generation result to align the generated relations with relations present in the local neighborhood of the query entity in the knowledge graph. GMT-KBQA [51] first retrieves relevant entities and relations through bert-based deep retrieval, then concatenates the information for T5 generation. To improve the retrieval result, it leverages cross-encoder to rerank the candidates, and uses the same T5 structure to conduct relation classification and entity dismbiguation. RNG-KBQA [122] enumerates candidate logical forms in the knowledge graph, then ranks the candidates and concatenates them with query to form the prompt input to generate final logical form through a T5 model. Based on this idea, TIARA [123] also retrieve entity and schema besides logical forms, while a following work [322] retrieves top-k questions with BM25. Uni-Parser [134] retrieves relevant entities from knowledge graph using mention detection, cross-encoder ranker, and 2-hop paths extraction. It also considers enumerating tables and columns from databases. On obtaining the relevant information, it concatenates the top-k primitives with the query and generates logical forms through T5. BLLM augmentation [136] uses TIARA as the retrieval for relevant knowledge base elements, then performs in-context learning through black-box LLM such as GPT-4 for generating logical forms. ECBRF [135] follows the case-based reasoning paradigm [323], retrieving similar triplet with dense retriever and constructing prompt input for BART or GPT-2 in-context learning. FC-KBQA [324] extracts relevant class, relation, and entity given a question. For class and relation, it uses BM25 as retriever and a Bert-based cross-encoder as re-ranker. For entity, it follows the mention detection paradigm. To compose all the component candidates, it applies T5 model to generate logical expression. StructGPT [325] extracts relevant information, including triplets and nearest entities, to form prompt for subsequent LLM. KAPING [326] builds prompt including the user query and the retrieved relevant facts (through entity matching), then generates the answer through LLM. Another work [327] also follows the retrieve-then-generate paradigm, replacing the retrieval with a relation distribution generation model for weighted triplets. Retrieve-Rewrite-Answer [328] first retrieves subgraph using hop prediction, relation path prediction, and triplet sampling. It then performs KG-to-text and zero-shot generation with retrieved subgraphs as prompt. Keqing [329] first decomposes a complext question into simple sub-questions through finetuned LLM, then retrieves similar sub-question template by dense retriever RoBERTa to extract candidate entities from knowledge graph, and finally generates answer through ChatGPT given relevant entities as context input. To probe the deep understanding of natural language in LLMs with formal languages, a research work [330] explores the capability of formal language understanding and formal language generation. It leverages retrieved pairs to perform in-context learning. For understanding, it uses tree edit distance to retrieve similar logical forms, while for generation, it uses BM25 to retrieve similar natural language queries. Interactive-KBQA [331] treats LLM as agent and KG as environment. In each step, the LLM conducts entity-linking and one-hop retrieval on KG, then generates current thought and action, until obtaining the final answer.
Latent representation-based RAG is also employed in KBQA. ReTraCk [332] links entities using mention detection, and retrieves schema items using dense retriever Bert. It then generates logical forms by LSTM, incorporating retrieved items through knowledge-specific rules. SKP [146] concatenates triplets with the query and uses fusion-in-decoder technique in inference. It adds a pretraining stage with a knowledge-aware MLM loss on triplets and knowledge constrastive loss with respect to the retrieved items. DECAF [145] forms Resource Description Format knowledge base triplets into sentences, then concatenates sentences with the same head entity into documents. It retrieves relevant documents using BM25 sparse retrieval or Bert dense retrieval, then leverages Fusion-in-Decoder technique to generate logical form and direct answer given each (query, document) pair as input. It combines the output to obtain the final answer. KD-CoT [147] uses the same dense retriever and fusion-in-decoder generator as DECAF. It follows a Chain-of-Thought paradigm, iteratively performing retrieval, generation, and verification until the CoT is finished.
IV-G2 Knowledge-augmented Open-domain Question Answering
Structured knowledge is often leveraged for open-domain question answering to achieve outstanding performance.
The use of latent representation-based RAG, particularly the fusion-in-decoder technique, is prevalent for knowledge-augmented open-domain question answering. UniK-QA [144] concatenates the text forms of the components in a triplet and build document pool for retrieval. For a given question, it leverages dense retriever for relevant documents, then performs fusion-in-decoder technique to incorporate the information for answer generation.
Figure 15: Architecture of KG-FiD [333] model.
KG-FiD [333] conducts the retrieval and generation as normal FiD. It add re-ranking in two ways: the first is to use a graph attention network on the graph formed by retrieved documents; the second is to use the hidden states in the generator. OREOLM [334] empowers LLM with knowledge reasoning paths. Concretely, it uses entity linking to determine the initial state, then conducts contextualized random walk on KG to get reasoning paths, whose entity value memory are combined into the hidden states of LLM for better generation. GRAPE [335] constructs bipartite graph for each pair of question and retrieved passage, then builds bipartite graph on entity for fusing knowledge. It leverages FiD as backbone model and generate answers. SKURG [336] forms a knowledge graph using text and image data, then updates each source’s representation with their relevant sources. It then uses a specially designed decoder to iteratively retrieve and generate. It conducts cross-attention with all the sources, then retrieves the source with highest score and concatenates to the original input embedding; if a gate score does not exceed the threshold, it starts to generate the real answer, otherwise the retrieval stage re-starts.
With the rapid development of LLMs, query-based RAG is emerging as a new standard. DIVKNOWQA [337] develops a retrieval-augmentation tool, including sparse retrieval on structured knowledge, dense retrieval on texts, and sparql generation on KB. Through CoT-based LLM, it retrieves and re-ranks in each step, and generates the final answer. KnowledGPT [338] uses generated code to retrieve from both public and personal knowledge bases, so as to build prompt for LLM question answering. EFSUM [339] generates evidence-focused summary with retrieved relevant facts, then optimizes the summary to align the QA-specific preference with another generator and the filters for helpfulness and faithfulness. GenTKGQA [340] conducts subgraph retrieval through relation ranking and time mining, then employs GNN to incorporate structural and temporal information into virtual token representations for subsequent LLM generation. KnowledgeNavigator [341] analyzes complex questions and performs retrieval on knowledge graph through iterative filtering of relations with respect to core entities, so as to obtain relevant triplets. It then includes the triplets into prompt for generation.
IV-G3 Table for Question Answering
Tables, as another form of structured knowledge, also facilitates question answering.
Fusion-in-decoder style RAG is often used for table QA. Many retrieval-reader systems are proposed on the EfficientQA competition in NeurIPS 2020 [342]. The systems retrieve from texts or tables, then leverage extractive or generative models as readers. Dual Reader-Parser [343] retrieves relevant textual and tabular information using BM25, then re-ranks using cross-encoder Bert. It leverages Fusion-in-Decoder to combine the retrieved results and generate the final answer. Convinse [344] first conducts question understanding, then retrieves relevant information from heterogeneous resources (including knowledge base, tables, and texts) by entity linking then sparse retrieval. It finally uses fusion-in-decoder to incorporate retrieved evidence and generate answers. CORE [345] retrieves relevant tables and passages through dense representation, then re-ranks the retrieved results using query-generation score from a T0 linker. It uses FiD for final generation. RINK [346] follows the retriever-reranker-reader paradigm for table-augmented question answering. It designs a set-level reader-inherited reranker to get the relevance score of blocks (table segments). TAG-QA [347] retrieves from both tables and texts: for tables, it converts tables to graphs then performs GNN to select relevant contents; for texts, it uses BM25 for retrieval. It then leverages FiD for answer generation.
Tables can be incorporated into prompts for query-based RAG. T-RAG [348] retrieves relevant tables given a user query, then concatenates the query with the tables as a prompt to generate the answer through BART. OmniTab [349] conducts multi-task training to improve the performance of table question answering model. It retrieves relevant tables given a query, then concatenates them for masked-token pre-training. CARP [350] retrieves relevant tables and passages using entity linking, then extracts hybrid chain of retrieved knowledge, which is later used to construct prompt for generation. StructGPT [325] extracts relevant information from knowledge graph, table, or database, to form prompt and generate answers through LLMs. cTBLS [351] retrieves relevant tables through dense retrieval, then for each query, it forms prompt with ranked tables for answer generation. A recent work [352] first uses table-to-text techniques to integrate tabular data into corpora, then conducts experiments on both finetuning and RAG for question answering.
IV-G4 Others
Prototype-KRG [353] retrieves knowledge facts and dialogue prototypes, then integrates them into the GRU-based generator by both hidden states and logits. SURGE [354] retrieves relevant subgraphs using dense retrieval, then adds them into the input of the generator for dialogue generation. RHO [355] fuses KG embedding of relevant entities and relations into textual embeddings during the generation of open-domain dialogue. K-LaMP [356] retrieves entity in history queries to construct prompt for query suggestion. ReSKGC [148] linearizes all training triplets into text by concatenation, then retrieves relevant triplets using BM25, and generates completed triplet using T5 with fusion-in-decoder. G-Retriever [357] retrieves relevant nodes and edges from graph-based data, then constructs subgraph and performs graph prompt tuning for question answering based on textual graphs.
@@ IV-H RAG for Science
RAG has also emerged as a promising research direction for many interdisciplinary applications, such as strengthening molecular generation, guiding medical tasks and advancing computational research.
@@ IV-H1 Drug Discovery
The goal of drug discovery is to generate molecules that concurrently fulfill diverse properties.
Figure 16: Architecture of RetMol [54] model.
RetMol [54] incorporates a lightweight retrieval mechanism into a pre-trained encoder-decoder generative model. As shown in Fig. 16, this approach adopts the SMILES string [358] representation of molecules along with the ChemFormer model [359] , involving the retrieval and fusion of exemplar molecules with the input molecule. PromptDiff [360] proposes an innovative interaction-based retrieval-augmented 3D molecular diffusion model to facilitate target-aware molecule generation. PromptDiff retrieves a curated set of ligand references, i.e., those with desired properties such as high binding affinity, to steer the diffusion model towards synthesizing ligands that satisfy design criteria.
IV-H2 Biomedical Informatics Enhancement
Several recent studies have improved the expressiveness of LLM by retrieving information from biomedical domain-specific databases, thereby augmenting the model’s capabilities to provide valuable guidance for tasks in the medical field. PoET [361] is an autoregressive generative model based on a variant of Transformer that integrates a retrieval mechanism to enable prompt augmentation, thereby expediting the prediction of fitness properties for protein variants. Chat-Orthopedist [137] creates a retrieval-augmented ChatGPT using an external knowledge base on adolescent idiopathic scoliosis (AIS), enhancing LLMs with AIS-specific information to provide accurate responses to inquiries. BIOREADER [362] is the first retrieval-enhanced text-to-text transformer-based model for biomedical natural language processing, incorporating the retrieved literature evidence into the model using a chunked-cross attention mechanism. MedWriter [363] utilizes a hierarchical retrieval-augmented generation approach, where the Visual-Language Retrieval (VLR) module retrieves report-level templates, and the Language-Language Retrieval (LLR) module fetches sentence-level templates, to generate coherent and clinically accurate medical reports from images. QA-RAG [364], utilizing a dual-track RAG approach, leverages fine-tuned language model responses and user queries to accurately retrieve and synthesize regulatory guidelines, significantly improving task performance in pharmaceutical compliance.
IV-H3 Math Applications
Retrieval-augmented generation technology in mathematics streamlines problem-solving, boosts research innovation, and refines educational strategies. LeanDojo [365] enhances theorem proving with its ReProver model, utilizing retrieval-augmented techniques to efficiently select relevant premises from extensive mathematical libraries, thus improving automation and theorem generalization. RAG-for-math-QA [366] enhances math question-answering by leveraging retrieval-augmented generation to integrate content from a high-quality math textbook, aiming to improve the quality and educational relevance of LLM-generated responses for middle-school algebra and geometry queries.
@@ V Benchmark
CRUD-RAG [371] divides all RAG tasks into four categories, which are Create, Read, Update, and Delete, respectively, and also evaluates each category using text continuation, question answering (with single- and multi-document questions), hallucination modification, and open-domain multi-document summary. MIRAGE [372] is a benchmark designed for assessing the application of RAG in the medical domain, focusing on the comparison and optimization of medical question-answering systems’ performance.
@@ VI-B2 Efficient Deployment and Processing
Currently, several deployment solutions of query-based RAG for LLMs have been proposed, such as LangChain [379], LLAMA-Index [180], and PipeRAG [380]. However, for other types of RAG and generation tasks, there lacks a plug-and-play solution. In addition, given the extra overhead introduced by retrieval, and considering that the complexities of both the retriever and generator will continue to grow, achieving efficient processing in RAG remains a challenge, necessitating targeted system optimization.